### v1

# Building a Data Dashboard for Magic: The Gathering: Initial Design Considerations Preprint — arXiv

Abstract
This paper presents the initial stages of a design study aimed at developing a dashboard to visualize gameplay data of the Commander format from Magic: The Gathering. We conducted a user-task analysis to identify requirements for a data visualization dashboard tailored to the Commander format. Afterwards, we proposed a design for the dashboard leveraging visualizations to address players’ needs and pain points for typical data analysis tasks in the context domain. Then, we followed-up with a structured user test to evaluate players’ comprehension and preferences of data visualizations. Results show that players prioritize contextually relevant, outcome-driven metrics over peripheral ones, and that canonical charts like heatmaps and line charts support higher comprehension than complex ones such as scatterplots or icicle plots. Our findings also highlight the importance of localized views, user customization, and progressive disclosure, emphasizing that adaptability and contextual relevance are as essential as accuracy in effective dashboard design. Our study contributes practical design guidelines for data visualization in gaming contexts and highlights broader implications for engagement-driven dashboards.
{CCSXML}
<ccs2012> <concept> <concept_id>10003120.10003145.10003147.10010923</concept_id> <concept_desc>Human-centered computing Information visualization</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10003120.10003145.10011770</concept_id> <concept_desc>Human-centered computing Visualization design and evaluation methods</concept_desc> <concept_significance>500</concept_significance> </concept> </ccs2012>
[…]
1 Introduction
Recent research in modern board and card games focuses on leveraging data-driven methods, artificial intelligence (AI) models, and visualization techniques to support decision-making, strategy development, and player understanding. Among these games, Magic: The Gathering (MTG) stands as the progenitor of the wider trading card game (TCG) genre as it is known today and the world’s largest tabletop collectible card game [ohm2023magic, zhu2023intelligence]. Specifically for MTG, some studies focus on generalized card representations, where multi-modal embeddings combine textual, numeric, visual, and usage-based features to capture card properties across expansions [bertram2024learning]. Other researchers tackle how game complexity has evolved through the years [magruder2022conservative] or card synergies [xue2025cardsynergy] using data visualizations. Another major area of research examines how AI agents and prediction models help drafting and human decision-making [ward2021ai, bertram2021predicting, zhu2023intelligence, alvin2021toward]. Traditional drafting is a complex process requiring players to balance immediate card value against long-term deck coherence. Beyond mechanics and data, perceptual and aesthetic factors also play a significant role. For instance, Kallabis et al. [kallabis2025deceptive] showed that visual styling can bias how players perceive card strength, even when objective properties remain constant. Finally, other studies focus on the gamer culture of MTG [trammell2010magic, falcao2021conservatism], predicting personality traits and color preferences [cui2025identity], or the cultural aspects of designing cards [fornazari2016gender, zanescu2022designing].
Together, these studies highlight an emerging research space at the intersection of AI, visualization, and trading card games. However, we identified two critical research gaps in MTG related research. First, there is a lack of research on MTG’s most popular format, Commander. This format has become especially popular due to its large deck size, singleton rules, and focus on synergy-driven gameplay. This richness makes deck construction and strategy development both rewarding and challenging, as players must navigate vast card pools, hidden synergies, and shifting metagames. Second, although community-driven resources and APIs (e.g., Scryfall (www.scryfall.com) and MTGJSON (www.mtgjson.com)) and deck-building websites (e.g., Archidekt (www.archidekt.com) or Moxfield (www.moxfield.com)) provide comprehensive card data or data visualizations to track deck metrics, there is currently no research on visual analytics for MTG gameplay data, leaving players without interactive tools to explore strategic patterns or deck performance.
In this paper, we report on the initial stages of a design study to develop a visualization dashboard for Commander. Our approach draws from design studies [sedlmair2012design] in visualization research, focusing on supporting players in understanding synergies, situating their decks within broader strategic contexts, and assessing their performance against other players. Specifically, we contribute with:
  • A user-task analysis to identify requirements for a data visualization dashboard tailored to the Commander format.
  • An evaluation of users’ comprehension and preferences through a structured user test of data visualizations.
2 Related Work
Within human-computer interaction (HCI) research, user-centered design [norman1986user] has become a well-established framework for developing systems that prioritize usability and user acceptance [fdis20099241]. This approach emphasizes integrating the needs, skills, and preferences of stakeholders into the design process [kopec2017livinglab, wilkinson2014applying]. Engaging users not only accelerates prototyping cycles [signoretti2015trip, ferreira2014elderly], but also enhances the overall quality of outcomes, often leading to more impactful and widely adopted solutions [lindsay2012engaging]. A specific branch of user-centered design, participatory design (PD), has been employed to uncover relevant requirements in case studies [frishberg2011interactive, kahng2017cti]. By fostering co-creation, it enables users to actively contribute to the identification of suitable design directions[lee2017steps]. Evidence shows that PD increases customer satisfaction and loyalty in private-sector contexts [grissemann2012customer], while in the public sector it has become central to driving social innovation [voorberg2015systematic].
2.1 Design Studies in Information Visualization
Design studies can be considered a methodological extension of PD, as they frequently involve users throughout the design process [lam2017bridging]. Sedlmair et al. [sedlmair2012design] define a design study as “a project in which visualization researchers analyze a specific real-world problem faced by domain experts, design a visualization system that supports solving this problem, validate the design, and reflect about lessons learned to refine visualization design guidelines.” This problem-driven methodology has been applied across domains (e.g., [goodwin2016constraint, kerzner2017graffinity, nobre2018lineage]) to capture existing practices and improve visualization solutions [leon2020co]. While the approach entails certain challenges and pitfalls [sedlmair2012design, kerzner2018framework], close collaboration between researchers and stakeholders has been shown to strengthen trust, rapport, and co-ownership [goodwin2016constraint]. Moreover, this collaboration enables stakeholders to bring in their domain expertise and jointly explore visualization opportunities [goodwin2016constraint].
Recent work in visualization and interactive systems demonstrates the value of user-centered, multi-view, and explainable approaches for complex data domains (e.g., [wang2022extending, ruan2024violet, kandel2024pd, sarigai2025dciwebmapper]). dciWebMapper [sarigai2025dciwebmapper] provides a scalable framework for exploring high-dimensional geospatial data through coordinated, interactive views, illustrating how linked visualizations support pattern discovery. In healthcare, PD-Insighter [kandel2024pd] shows that participatory design can enable clinicians and patients to monitor daily behaviors effectively through temporal and spatial visualizations. For AI and complex model interpretation, Violet [ruan2024violet] visualizes quantum neural network activations to make opaque decision processes understandable, while Wang et al. [wang2022extending] apply user-centric explainable AI design to graph neural network outputs in drug repurposing, emphasizing domain-specific interpretability. A small subset of works focuses on investigating practical knowledge, identifying user challenges and visualization opportunities rather than prototype implementation to support comprehension and navigation within other domains (e.g., [li2023knowledge, wolf2025identifying]). Together, these studies highlight design principles that are directly applicable to building dashboards for complex domains such as trading card games: integrating multi-view coordination, supporting exploratory analysis, explaining complex relationships, and grounding the interface in user needs and domain expertise. Our work builds upon these findings to identify user requirements and propose data visualizations that directly support MTG analysis.
2.2 Data Visualizations for Magic: The Gathering
State-of-the-art research offers limited evidence of applying data visualization to explore card synergies, gameplay data, or card collections. Nevertheless, several online resources provide visual and analytical tools for MTG, ranging from community-driven dashboards to professional-grade platforms. For example, Vincenzo Galante’s Looker dashboard (www.github.com/VincenzoGalante/magic-the-gathering) presents a comprehensive analysis of MTG data through common visualization types, including horizontal bar charts for yearly card releases and pie charts for color distribution (Figure 1). Similarly, another exploratory analysis (www.colinberan.com/2023/06/11/exploratory-data-analysis-of-magic-the-gathering-cards) employs pie charts for color balance and vertical bar charts for card rarity. Online deck-building platforms adopt comparable visual conventions. For instance, Archidekt uses pie charts to represent color counts and bar charts to illustrate mana-value distributions (Figure 2).
Figure 1: Screenshot of the Galante’s dashboard.
Figure 2: Screenshot of Archidekt’s data visualizations for collections.
We found two resources that are aligned our objectives. First, Pacheco’s Commander dashboard (www.github.com/gabi-pacheco/MTG-Commander-Analysis) analyzes 53 Commander matches played between four players from June 2023 to August 2024 (Figure 3). This dataset captures deck usage, player performance, win conditions, and match outcomes, offering insights into player behavior and strategic tendencies. The dashboard also integrates a variety of visual forms (such as arc diagrams, area charts, and scatterplots) that can be interactively filtered by decks and players. A major limitation is that this project is custom made and does not allow other players to view their data. It also lacks interactive features and the ability to gather insights that are not explicit in the visualizations. Second, Playgroup.gg is a platform for tracking MTG Commander games, designed to capture game details and provide deep statistical insights for both casual and competitive players through extensive data visualization features. Key visualization features include (i) game and deck stats dashboards with charts for metrics such as win rates, average turn lengths, commander usage, power level estimates, and leaderboard positions and (ii) individual player dashboards featuring most-played decks, historical performance, and social stats such as rivalries and achievements (Figure 4). Although this project showcases the highest data diversity, the charts are spread across the website without an easy way to integrate them or allow for filtering. Color palettes for categorical encodings also hinder users’ ability to identify individual marks. Another reported limitation hinges on how challenging it is to find custom and granular stats visualizations.
(a) Overall.
(b) Deck performance.
(c) Deck usage consistency.
Figure 3: Screenshots of the dashboard developed by Pacheco.
(a) Deck dashboard.
(b) Player dashboard.
(c) Playgroup dashboard.
Figure 4: Screenshots of the dashboards developed by Playgroup.gg.
Together, these resources provide valuable reference points for developing our own visualization dashboard for Commander, highlighting the most common tasks, data dimensions, and chart types employed in existing practice. However, their limitations also reveal key opportunities for improvement. First, most existing dashboards focus on descriptive statistics and isolated charts rather than on facilitating comparative or exploratory analysis. This restricts users’ ability to identify deeper gameplay patterns such as synergies between commanders, temporal trends in performance, or contextual factors influencing win conditions. Second, visual encodings often rely on color as the primary differentiator, leading to accessibility issues and difficulties in distinguishing overlapping marks. Third, limited interactivity and lack of customization prevent users from tailoring visualizations to their own analytical goals or filtering data according to personal playgroup contexts. Finally, the fragmentation of visual components across different pages impedes holistic understanding and sensemaking. In the next section, we present our dashboard design methodology focused on addressing these issues while catering for player needs.
3 Methodology
Sedlmair et al.’s Nine-Stage Framework [sedlmair2012design] provides a structured methodology for conducting design studies in visualization research and is widely recognized for helping researchers address real-world problems in close collaboration with domain experts. The stages are as follows: (i) learn: researchers gain in-depth knowledge about the target domain and context, primarily through close engagement with domain experts; (ii) winnow: the initial set of possible problems is narrowed down to the most important and feasible ones for focused study; (iii) cast: researchers frame and abstract the chosen problem, identifying main data, tasks, and requirements that should guide the design; (iv) discover: exploration of possible design solutions is conducted, often through brainstorming, prototyping, and iterative feedback; (v) design: selected visualization solutions are refined and finalized based on insights gained from earlier exploration and input from domain experts; (vi) implement: the visualization system or tool is developed to realize the chosen design, paying attention to usability and technical constraints; (vii) deploy: the system is delivered to domain experts for real-world use and feedback, often including iterative refinement as issues arise; (viii) reflect: researchers engage in analysis and reflection about what worked, what did not, and lessons learned for future guidance; and (ix) write: the process and findings are documented and reported for dissemination to the broader research community. The framework emphasizes the iterative and overlapping nature of these stages rather than strict linearity.
In our study, we followed Sedlmair et al.’s structured approach to establishing effective collaborations. In particular, we went through stages from learn to design with a focus on understanding players’ needs and pain points regarding current data collection and analysis tools as well as which data visualizations may be adequate in this context to provide untapped insights of Commander gameplay. Throughout the process, we balanced usability with analytical depth, ensuring that the resulting dashboard provided actionable, context-sensitive insights while remaining approachable for players with varying levels of engagement, technical skill, and visualization literacy. In turn, this study forms the foundation for the subsequent design of the visualization dashboard. Experimental materials, code, and results are available at an online repository (https://osf.io/rsz9g/overview?view_only=e50cc8aff82e47c580bb422633599487).
4 Precondition Phase
This section presents the stages of learn, winnow, and cast.
4.1 Learn
As we mentioned, we focus on Commander (https://magic.wizards.com/en/formats/commander) (formerly known as Elder Dragon Highlander or EDH), one of the most popular and distinctive formats in MTG. Unlike competitive one-on-one formats, Commander emphasizes creativity, social interaction, and long-term strategy, making it particularly appealing to casual and community-oriented players. Each deck contains exactly 100 cards, and with the exception of basic lands, no duplicates are allowed. Central to the format is the choice of a Commander, which is usually a legendary creature or planeswalker, that defines the deck’s color identity through its mana symbols and shapes its overall strategy. Games are most often played with four players, each starting at 40 life points, which is double the usual starting total. The Commander itself is always available to cast from the Command Zone and can be replayed with an increasing two-mana tax each time it is removed. In addition to regular win conditions, the format introduces a unique rule called Commander Damage, where a player loses the game if they receive 21 combat damage from a single Commander over the course of the game. Commander is strongly defined by its synergy-driven deck-building. The singleton rule and the enormous card pool of more than 20,000 unique cards foster highly diverse and often thematic decks. Social dynamics, such as table politics, threat perception, and temporary alliances, further distinguish the format from competitive play. This combination of diversity, strategic depth, and strong community engagement has made Commander a cornerstone of modern MTG.
4.2 Winnow
After identifying the problem domain, we continued the design study with the winnow stage. During this phase, we engaged in initial discussions to identify potential collaborators within the MTG community who had a strong alignment between their domain needs and our research objectives. The Commander format helped raise an ecosystem of deck-building platforms, online discussions, and data-driven tools, making it a particularly rich case study for research in visualization and analytics. Among these tools, Gauntlet (www.gauntletapp.com), Playgroup.gg (www.playgroup.gg), Dragon Counter (www.dragoncounter.com), LifeLinked (www.sites.google.com/view/lifelinked/home), MTG Life Counter: Lotus (www.lifecounter.app) and Mythic Track (www.mythictrack.com) stand as the preferred tools to track track life, wins, or more general gameplay history. We analyzed these tools to understand the main features, data availability, and the potential to host a dashboard for Commander data. Our analysis shows that most tools do a satisfactory job of tracking generic gameplay data (e.g., damage, game record-keeping), but they either track life totals or game records. Among them, Playgroup.gg stands out as the exception, being able to track life and other gameplay data seemingly. Playgroup.gg also offers a rich history of game statistics and static dashboards with data visualizations, allowing players to explore their data and compare it against other decks, players, and playgroups (Figure 4). Therefore, we contacted the founders and developers of Playgroup.gg, Maran and Johannes, to establish a collaboration.
4.3 Cast
The subsequent cast stage involved formalizing this collaboration, confirming mutual expectations, and committing resources to jointly develop the visualization dashboard. We identified the front-line analysts as every Playgroup.gg user since they interact directly with gameplay data and would use the resulting dashboard. Additionally, the Playgroup.gg developers are the gatekeepers, who hold authority over data access and project approval. Finally, we, the authors, act as HCI researchers and translators since we need to (i) abstract the domain problems into a more generic form considering larger-context domain goals, (ii) conduct a user task analysis, and (iii) design and evaluate data visualizations to satisfy user needs related to the domain problems. Establishing these roles early ensured that we could proceed efficiently and conduct a successful iterative design of the Commander dashboard.
5 Core Phase
This section presents the stages of discover and design.
5.1 Discover
During the discover stage, we focused on understanding which data and insights players were most interested in regarding the Commander format, in order to identify opportunities for visualization.
5.1.1 User-Task Analysis
We conducted a user-task analysis using a structured questionnaire designed to capture both the strengths of Playgroup.gg’s current practices and the challenges users encountered when exploring its existing static dashboards. The questionnaire was organized into four main sections to assess users’ interests and needs concerning: (i) gameplay data, (ii) relationships between different attributes, (iii) visual analytical tasks, and (iv) data comparison contexts (e.g., comparisons between a player’s decks, within a specific playgroup, or across all users). Each item was rated on a 5-point Likert scale, ranging from very uninteresting (-2) to very interesting (2). We also collected demographic data to understand our sample and assess its diversity.
5.1.2 Results
The questionnaire was released on Reddit and Discord servers focused on MTG as well as on Playgroup.gg’s webpage. A total of 68 participants filled in the questionnaire, with 58 (85.29%) currently being Playgroup.gg users. Players showed the highest interest in commonly used performance statistics, with win rate (M = 1.62, SD = 0.69) and wins versus losses (M = 1.57, SD = 0.70) rated as most important, followed by deck and player rankings (M = 1.40, SD = 0.88). Metrics related to gameplay events, such as damage and healing (M = 1.04, SD = 1.16), number of rounds (M = 0.99, SD = 0.97), and starting player position (M = 0.87, SD = 1.13), were moderately interesting, whereas less common statistics like estimated deck power level (M = 0.82, SD = 1.23), kill frequency (M = 0.76, SD = 0.92), and ELO (M = 0.71, SD = 1.21) received lower interest. ELO is a method for calculating the relative skill levels of players in zero-sum games. Players expressed little to negative interest in peripheral or largely chance-driven metrics such as the number of priority changes (M = -0.13, SD = 1.21), number of dice rolls (M = -0.38, SD = 1.21), and pause duration (M = -0.63, SD = 1.08). Overall, these results suggest that users prioritize metrics directly related to game outcomes and player/deck performance over auxiliary or stochastic elements of gameplay.
Players expressed the greatest interest in relationships that directly link game outcomes to key gameplay variables, with the strongest preferences for win rate by commander (M = 1.71, SD = 0.62), wins by commander (M = 1.49, SD = 0.89), and deck/player rank by commander (M = 1.43, SD = 0.74). Other moderately interesting relationships included win rate by color identity (M = 1.28, SD = 1.05), win rate by starting position (M = 1.22, SD = 1.03), and kill frequency by commander (M = 1.21, SD = 1.13). Relationships involving less central metrics, such as damage/healing by commander or turn, deck power level by starting position, and deck/player rank by game duration, received moderate to low interest (M $\approx$ 0.5–0.9). Similar to the previous analysis, players showed little to negative interest (M $\leq$ 0) in relationships that included peripheral factors, including game priority changes, pause durations, and salt score. “Salt score” is an informal or playful metric used to measure a player’s level of frustration, annoyance, or “tilt” during gameplay. Overall, these findings indicate that users prioritize analyses that reveal how specific commanders, colors, and player performance metrics influence game outcomes, while relationships involving marginal or indirect variables are of limited interest.
Players showed the greatest interest in dashboards that provide a Playgroup view, focusing on members of a specific playgroup (M = 1.32, SD = 0.95), followed by a User view, which isolates data from the individual player’s decks (M = 1.07, SD = 1.20). The Global view, aggregating data across all playgroups, was considerably less interesting to users (M = 0.21, SD = 1.31), suggesting that players value contextually relevant comparisons over broad, population-level analyses. These values are in line with the view of Commander as a casual and social format. Regarding insights, users prioritized analyses that reveal strategically actionable patterns, with the highest interest in exploring the relationship between games played, games won, and the number of kills per commander (M = 1.22, SD = 1.08). Other moderately interesting insights included the popularity of color identities and commanders (M = 0.97, SD = 1.09), while metrics such as time spent per turn, damage per turn by commander, and damage type distributions per commander received lower interest (M = 0.46–0.71).
Afterwards, we analyzed the qualitative data, where players suggested a wide range of additional metrics, analyses, and insights that could enhance the visualization of Commander gameplay. Several ideas focused on deck- or player-level performance, including metrics like archenemy tracking to measure which players or decks are perceived as threats, average turn to cast a commander, or commander casts per game. Other gameplay aspects included identifying trends in personal skill and deck effectiveness. Regarding analysis types, players expressed interest in metrics that could reveal actionable patterns such as the ability to compare performance across different playgroup sizes or against players with differing skill levels. Overall, these results indicate that players value strategically actionable, turn-sensitive, and contextually relevant metrics, with a preference for views and insights that emphasize comparisons within their own playgroup or deck context. They also seek the flexibility to track long-term trends, deck evolution, and analyses that connect directly to game outcomes and strategic performance.
5.2 Design
In the design stage, we translated the insights gathered during the discover phase into concrete visualization solutions for Playgroup.gg. This process involved generating and validating data abstractions, visual encodings, and interaction mechanisms that could effectively communicate gameplay and strategic information to players. Before narrowing the proposal space in consultation with the Playgroup.gg founders, we iteratively explored multiple design alternatives, including different chart types, filtering options, and comparison contexts (e.g., individual decks, playgroup-wide, or global views). We ran two rapid prototyping workshops to create sketches and mockups of an interactive dashboard to enable quick feedback loops, allowing us to refine visual encodings and interaction features based on domain expert input.
5.2.1 The Dashboard
Figure 5 presents the dashboard mockup the authors proposed to the Playgroup.gg team. The dashboard provides a comprehensive overview of player and deck performance in the Commander format, integrating multiple coordinated visualizations to support analysis at both granular and aggregate levels. The top section features a parallel coordinates plot summarizing key metrics such as number of wins and games played, damage, healing, ELO, and estimated deck power, enabling users to quickly compare multiple commanders across these dimensions. To the right, bar charts illustrate the frequency of specific game end states, facilitating the identification of deck strategies. Below, a scatterplot maps games played against games won, with shading encoding average kills, providing a clear view of performance trends, while an adjacent line chart tracks damage dealt per turn by commanders, supporting temporal analysis. The lower-left heatmap shows seating positions across multiple metrics, allowing rapid assessment of relative performance by play order. Finally, an icicle plot at the bottom-right encodes the number of decks per color identities and number of colors with an extra brightness encoding for player “saltiness”, highlighting patterns in player experiences and specific color combinations and commanders. Finally, the dashboard includes panels for active filters and customization instructions, ensuring that users can tailor views to specific analytical questions and comparative contexts.
Figure 5: Mockup of the proposed dashboard for Commander data analysis for Playgroup.gg.
5.2.2 Evaluation
We then met with the Playgroup.gg team to gather their feedback on the proposed dashboard. While they praised the overall design, they noted one weakness: using only a black stroke for the marks in the line chart made it difficult to track a single line across the x-axis. After consideration, we decided to retain the black stroke, as the goal of the visualization was not to individually identify commanders since this functionality was already supported through filtering in the right-hand panel. Instead, we want to analyze the overall evolution of damage per turn. Following this, we collaborated with the Playgroup.gg team to develop a questionnaire designed to assess whether users could understand and extract insights from the charts. The questionnaire contained one section per chart, with each item structured as a multiple-choice question with only one correct answer. Each item required participants to interpret a chart and respond accordingly. To reduce random guessing, participants were explicitly instructed not to guess and were given the option to select I don’t know if undecided. We framed the questionnaire as a quiz, awarding one point per correct answer, to foster engagement. Finally, participants were invited to provide open-ended feedback as well as demographic information to contextualize the sample.
5.2.3 Results
Similar to the previous questionnaire, we released the new one on Playgroup.gg’s website, Reddit, and Discord servers. A total of 83 participants, from which 77 (92.77%) were already Playgroup.gg users, answered the 33-item questionnaire. Participants demonstrated relatively strong comprehension of the dashboard, with an average raw score of 25.05 (SD = 6.22) out of 33. The distribution of scores ranged from 6 to 33, indicating some variability in performance. The 95% confidence interval for the mean raw score [23.69, 26.41] suggests that, on average, participants were able to correctly answer around three-quarters of the items. Corrected scores, which penalize guessing, produced a similar pattern with a mean of 23.97 (SD = 6.59) and a range of 4.85 to 33, further supporting that most users could extract accurate insights from the visualizations.
When broken down by chart type, performance varied substantially. Heatmaps yielded the highest mean accuracy at 98.19%, with nearly all participants correctly identifying seating position outcomes. Similarly, parallel coordinates plot showed high comprehension at 91.57%, indicating that users could effectively extract relative comparisons. The line chart also performed strongly, with a mean accuracy of 78.92%, suggesting that participants were able to interpret temporal patterns of damage per turn despite earlier concerns about line distinguishability. Bar charts showed a mean accuracy of 76.81%. While participants consistently identified dominant categories (100% accuracy for the most frequent win condition), comparative reasoning (e.g., estimation of proportional differences between conditions) was more error-prone, dropping to just over 50% accuracy. This highlights that while categorical comparisons are broadly understood, fine-grained quantitative estimation remains more difficult.
Scatterplots, while central to the dashboard, proved more challenging, with an overall mean accuracy of 67.47%. Although participants easily identified high-level features such as the most played commander (80% accuracy) or the commander with 62 games (87%), questions requiring precise interpretation of winrates or cross-referencing multiple attributes (e.g., kills at a specific winrate) saw lower performance, with some items falling below 50%. This suggests that scatterplots supported broad pattern recognition but demanded higher cognitive effort for precise lookups. The icicle plot showed a similar trend, with an overall mean accuracy of 75.30%. Participants were able to identify structural features such as color groupings and relative saltiest categories (above 75% accuracy for most items), but struggled with items requiring pinpoint identification of individual commanders, where accuracy dropped to 49%.
Regarding participant feedback, the most consistent theme across responses was the difficulty of interpreting the dashboard in black-and-white. Participants repeatedly emphasized that “colors would make these much easier to read” and that “it was so much harder to figure out in black and white than it would be with color.” This concern was especially salient for line charts, where overlapping strokes made it “virtually impossible to track the line graphs when they intersect.” Several respondents also requested colorblind-friendly palettes or user control over colors. Participants also reacted to charts individually, citing the heatmap and bar charts as “clear” and “my favourite visualization of data.” Scatterplots drew mixed reactions. While some appreciated their potential, many found the use of shading to encode kills unintuitive (“might be better with number of kills as size of the circle vs the shade”). Line Charts were seen as useful in theory but confusing in practice, particularly without color (“line graphs are far too confusing… color will help”). Icicle Plots received the harshest criticism, being described as “an eye-sore,” “horrible readability,” and even “gave me a headache.” A few users suggested mitigations like heavier boundaries or interactive callouts, but many implied they would avoid this visualization altogether.
Many participants requested interactive features, such as hovering or clicking to reveal precise values, filtering by attributes (e.g., power level, game length), and brushing/selection across linked charts. Customization was also a recurring theme. Users wanted the ability to “remove/add charts at your will,” “change each graph to show different states,” or select their preferred chart type for a given metric. This reflects a desire for dashboards that adapt to individual preferences and reduce information overload. Several participants also noted that the dashboard felt “too information dense” and “intimidating” when all charts were displayed at once. They suggested breaking the dashboard into sections, displaying charts one at a time, or providing quick-glance summaries (e.g., “top 3 winrates” or “saltiest commanders”) before drilling into detail. Additionally, participants mentioned the need for clearer separation between charts (“hard edges to separate each chart”) and consistency in elements like legend placement.
Finally, respondents proposed additional data or visual enhancements that would improve interpretability and engagement: (i) using commander images, mana symbols, or icons for easier recognition, (ii) providing labels or percentages directly on bars and data points, (iii) adding missing metrics such as average turn time or power level filters, and (iv) offering light/dark modes for accessibility and comfort. Despite criticisms, many participants expressed enthusiasm about the concept. The dashboard was called “very interesting,” “a good add in the Playgroup app,” and “great information displayed.” A few noted that the charts would be “useful after getting familiar with them,” highlighting the potential learning curve but also the value of the information once mastered.
5.3 Reflect
The results of this study allow us to confirm, refine, and propose design guidelines for dashboards supporting gameplay analysis in digital platforms such as Playgroup.gg. First, our findings confirm existing guidelines emphasizing the importance of contextually relevant, outcome-driven metrics: players consistently valued statistics like win rate and wins versus losses, while peripheral or stochastic measures such as dice rolls or pause duration were rejected as uninformative. Second, we refine guidelines for visualization choice by showing that although canonical charts such as heatmaps, parallel coordinates, and line charts supported high comprehension (M $\geq$ 78.9%), scatterplots and icicle plots posed substantial interpretation challenges, particularly under non-ideal presentation conditions such as black-and-white rendering. These results suggest the need to refine design guidance around complex or unconventional charts, recommending their use only when paired with interactivity (e.g., hover-to-reveal, filtering, brushing) or simplified encodings. Third, our findings reject the assumption that global or population-level dashboards are inherently useful: participants strongly preferred playgroup- and user-level views, suggesting that more aggregated, population-level data provides broader insight and is therefore more valuable. Finally, we propose new guidelines emphasizing user customization and progressive disclosure. Participants frequently requested the ability to tailor which charts were shown, select preferred encodings, and control dashboard density (e.g., quick summaries before detail), underscoring that adaptability is as critical as accuracy in supporting diverse player needs. Together, these reflections extend design study knowledge by grounding visualization guidelines in the practices and expectations of a specific yet illustrative community, while pointing to generalizable principles for balancing interpretability, customization, and contextual relevance in dashboard design.
Future work should review the proposed design in light of our results and continue with a prototype implementation and deployment in the Playgroup.gg website. Additionally, researchers may further investigate how interactive features can mitigate the challenges observed with complex visualizations such as scatterplots and icicle plots, particularly in mobile or constrained environments where overlapping elements are harder to trace. Further studies could explore adaptive dashboards that personalize chart selection and level of detail to different user profiles, such as novice players seeking simple summaries versus expert players interested in advanced cross-metric analyses. Incorporating player-centered enhancements such as commander images, mana symbols, or icons may improve recognition and engagement, and systematic testing of accessibility features (e.g., colorblind palettes, light/dark modes) could yield generalizable guidelines. Importantly, future research should broaden the participant base beyond Playgroup.gg users. Since our sample primarily consisted of individuals already interested in statistics and data tracking, their preferences may not fully represent the general Commander audience. Comparing results across both data-inclined and casual player populations would clarify whether the observed prioritization of strategic, outcome-driven metrics reflects a universal trend or a community-specific bias. Finally, longitudinal research is needed to understand how player preferences and comprehension evolve over time as familiarity with the dashboard grows, helping to refine guidelines around learning curves and long-term adoption.
6 Conclusions
This design study investigated the design and evaluation of a dashboard for analyzing gameplay data in MTG Commander, implemented in collaboration with the Playgroup.gg platform. Through iterative design, user feedback, and large-scale questionnaire testing, we examined which metrics, analyses, and visualization types players found most valuable, and how effectively they could interpret the resulting dashboard. Overall, this study contributes empirical evidence and practical guidelines for designing dashboards in gameplay analysis contexts, while also informing broader visualization research on user-centered design, comprehension, and engagement.
Acknowledgments
Work supported by national funds through Fundação para a Ciência e a Tecnologia, I.P. (FCT) under projects UID/50021/2025, UID/PRR/50021/2025, and UID/315/2025.